# Bonsai: Gradient-free Graph Distillation for Node Classification

### v1

@@ B.4 Additional Experiments
Cora 0.87 0.05 0.01 1.68 2.61
Citeseer 0.97 0.07 0.02 1.71 2.77
Pubmed 8.89 1.39 0.64 13.93 24.85
Reddit 682.80 293.15 275.52 193.99 1445.45
@@ B.5 Additional Training Times
Figure 8: Comparing time required to train models on datasets distilled by GCond, Gdem, Bonsai, and Gcsr at 0.5% of Cora, Citeseer, Pubmed, and Reddit. Note that there is no distilled dataset for Reddit output by Gcsr due to OOT.
Figure 9: Comparing time required to train models on datasets distilled by GCond, Gdem, Bonsai, and Gcsr at 1% of Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit. Note that there is no distilled dataset for Flickr, Ogbn-arxiv, Reddit output by Gcsr due to OOT.
Figure 10: Comparing time required to train models on datasets distilled by GCond, Gdem, Bonsai, and Gcsr at 3% of Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit. Note that there is no distilled dataset for Flickr, Ogbn-arxiv, Reddit output by Gcsr due to OOT; and distilled datasets produced by GCond and Gdem for Reddit cannot train on standard 2-layered PyTorch Geometric Gat and Gin due to OOM.
Fig. 8 presents the training times on distilled datasets in Cora, Citeseer, Pubmed, and Reddit at 0.5% size. Figs. 9 and 10 presents the training times on distilled datasets in Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit at 1%, and 3% sizes. Bonsai is faster on all architectures. The gap between Bonsai and the baselines increase with increase in distilled dataset size. This is more pronounced in larger datasets like Flickr, Ogbn-arxiv, and Reddit because the number of edges is $O(\lvert V\rvert^{2})$ for the baselines and forward pass through Gnns is usually $O(\lvert E\lvert)$. Additionally, the distilled dataset created by GCond and Gdem for Reddit at 3% size cannot be used to train basic PyTorch Geometric models for Gat and Gin, resulting in out-of-memory (OOM) errors because the backward pass graph maintained by PyTorch is larger.

### v3

@@ Appendix B Experimental Setup
Across all datasets, except MAG240M, we maintain a train-validation-test split ratio of $60:20:20$. In MAG240M, we use a ratio of 80:10:10.
@@ B.4 Additional Experiments
Cora 0.87 0.05 0.005 1.68 2.60
Citeseer 0.97 0.07 0.005 1.71 2.75
Pubmed 8.89 1.39 0.030 13.93 24.24
Reddit 682.80 293.15 0.696 193.99 1170.64
@@ B.5 Additional Training Times
Figure 7: Comparing time required to train models on datasets condensed by GCond, Gdem, Bonsai, and Gcsr at 0.5% of Cora, Citeseer, Pubmed, and Reddit. Note that there is no condensed dataset for Reddit output by Gcsr due to OOT.
Figure 8: Comparing time required to train models on datasets condensed by GCond, Gdem, Bonsai, and Gcsr at 1% of Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit. Note that there is no condensed dataset for Flickr, Ogbn-arxiv, Reddit output by Gcsr due to OOT.
Figure 9: Comparing time required to train models on datasets condensed by GCond, Gdem, Bonsai, and Gcsr at 3% of Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit. Note that there is no condensed dataset for Flickr, Ogbn-arxiv, Reddit output by Gcsr due to OOT; and condensed datasets produced by GCond and Gdem for Reddit cannot train on standard 2-layered PyTorch Geometric Gat and Gin due to OOM.
Fig. 7 presents the training times on condensed datasets in Cora, Citeseer, Pubmed, and Reddit at 0.5% size. Figs. 8 and 9 presents the training times on condensed datasets in Cora, Citeseer, Pubmed, Flickr, Ogbn-arxiv, and Reddit at 1%, and 3% sizes. Bonsai is faster on all architectures. The gap between Bonsai and the baselines increase with increase in condensed dataset size. This is more pronounced in larger datasets like Flickr, Ogbn-arxiv, and Reddit because the number of edges is $O(\lvert V\rvert^{2})$ for the baselines and forward pass through Gnns is usually $O(\lvert E\lvert)$. Additionally, the condensed dataset created by GCond and Gdem for Reddit at 3% size cannot be used to train basic PyTorch Geometric models for Gat and Gin, resulting in out-of-memory (OOM) errors because the backward pass graph maintained by PyTorch is larger.
B.6 Carbon Emissions
Table 8: Estimated CO2 emissions from condensation of various methods in seconds at 0.5%. Emissions that are higher than training on the full dataset itself are highlighted in red. The least emission from condensation for each dataset is highlighted in green. CO2 emissions are computed as 10.8kg per 100 hours for Nvidia A100 GPU and 4.32kg per 100 hours for 10 CPUs of Intel Xeon Gold 6248 (Lacoste et al., 2019).
Dataset GCond Gdem Gcsr Exgc GC-Sntk Geom Bonsai Full
Cora 82.14 3.15 157.80 1.05 2.46 389.88 0.03 0.75
Citeseer 81.36 5.01 199.08 1.03 3.72 472.89 0.03 0.75
Pubmed 77.01 15.90 39.57 3.45 3.51 $\geq$540.00 0.3 1.53
Flickr 58.05 102.15 523.35 7.30 18.36 $\geq$540.00 1.42 5.40
Ogbn-arxiv 434.22 17.07 $\geq$540.00 46.49 366.54 $\geq$540.00 4.18 15.74
Reddit 903.36 602.94 $\geq$540.00 207.10 876.33 $\geq$540.00 17.34 72.77